OpenStudy (mikezack123):

Which system of equations is represented by the following graph? a graph of two intersecting lines on a coordinate plane. One passes through points (0, 3) and (6, 0). The other passes through points (0, −3) and (3, 0).

OpenStudy (jessiegonzales):

the images?

OpenStudy (mikezack123):

i dont have it... :(

OpenStudy (mikezack123):

@amistre64 @dan815 @beccaboo333

OpenStudy (mikezack123):

NTS:#34

OpenStudy (amistre64):

you simply need to define the lines that are defined by their respective points

OpenStudy (mikezack123):

@amistre64 how i do that

OpenStudy (mikezack123):

A. x − 2y = 6 x + y = 3 B. y = x + 3 y = −1 over 2x + 3 C. y = x − 3 y =1 over 2x + 3 D. x + 2y = 6 x − y = 3

OpenStudy (mikezack123):

THANKS FOR YOUR HELP EVERYONE ON THIS WEBSITE

OpenStudy (amistre64):

the process of determining the line between 2 points is covered long before you have to work out a system of 2 equations ...

OpenStudy (amistre64):

find the slope between a set of points and apply the point slope formula of a line to extablish the equation

OpenStudy (mikezack123):

i canceled out B and C

OpenStudy (mikezack123):

amistire